OpenStudy (anonymous):

you can estimate the diameter of a tree without boring through it by measuring its circumference. Solve the formula c= πd for d. Then find the diameter of an oak that has a circumference of 113 inches.

OpenStudy (unklerhaukus):

\[C=\pi d\]dividing by pi \[\frac C \pi=\frac{\pi d}\pi\]simplify the right hand side

OpenStudy (unklerhaukus):

\[\frac C \pi =\frac{\cancel \pi d}{\cancel\pi}\]\[\frac C \pi=d\] substitute the circumference into the equation \[C=113 [\text {in}]\] \[\frac {113 [\text {in}]} \pi=d\] \[?=d\]